Mathematics 241 – Multivariable Calculus Quiz 9 – December 10, 2010

Your Name:

A. (10) Set up (but do not evaluate) the double integral of f(x, y) = y^{2} over the quarter circle H = {(x, y, z) | x^{2} + y^{2} ≤ 25, x ≥ 0, y ≥ 0} using rectangular coordinates.

B. (10) Same function f(x, y) and H as in A, but set up the double integral using polar coordinates.